### v1

# Pretrained Hybrids with MAD Skills

Abstract
While Transformers underpin modern large language models (LMs), there is a growing list of alternative architectures with new capabilities, promises, and tradeoffs. This makes choosing the right LM architecture challenging. Recently-proposed hybrid architectures seek a best-of-all-worlds approach that reaps the benefits of all architectures. Hybrid design is difficult for two reasons: it requires manual expert-driven search, and new hybrids must be trained from scratch. We propose Manticore, [^2] The Manticore is a fearsome human/lion/scorpion hybrid from Persian mythology. a framework that addresses these challenges. Manticore automates the design of hybrid architectures while reusing pretrained models to create pretrained hybrids. Our approach augments ideas from differentiable Neural Architecture Search (NAS) by incorporating simple projectors that translate features between pretrained blocks from different architectures. We then fine-tune hybrids that combine pretrained models from different architecture families—such as the GPT series and Mamba—end-to-end. With Manticore, we enable LM selection without training multiple models, the construction of pretrained hybrids from existing pretrained models, and the ability to program pretrained hybrids to have certain capabilities. Manticore hybrids outperform existing manually-designed hybrids, achieve strong performance on Long Range Arena (LRA) tasks, and can improve on pretrained transformers and state space models.
1 Introduction
Transformers are the workhorse architecture for large language models and beyond, powering a vast collection of foundation models. While for years it appeared that the Transformers family would remain the undisputed standard, a recent Cambrian explosion of proposed architectures has taken place. Many of the new architectures achieve subquadratic complexity—in contrast to the quadratic complexity of self-attention in Transformers—by using local or linear attention [9, 6, 3, 45], or resurrecting and scaling recurrent networks [6, 9, 31], or by building on state-space modeling principles [13, 33, 32, 11, 14]. These approaches potentially promise to overturn the dominance of Transformers through more efficient training and inference.
However, no single new model is a clear overall winner when varying data modalities, tasks, and model sizes. Comparing architectures on a fixed task is fraught with difficulties [2]. Even if these issues are overcome, practitioners would have to experiment with and evaluate every architecture for each new task—an expensive proposition. Instead, seeking a best-of-all-worlds approach, researchers have proposed the use of hybrid models that mix multiple architectures. These hybrids, such as the MambaFormer [29]—a mix of the popular Mamba with a Transformer—have shown potential in maintaining the desirable properties of multiple model classes.
While promising, hybrids suffer from two main obstacles that stymie their adoption:
  • Manual Design. Hybrid architectures are hand-crafted, either by manually exploring the large search space of hybrids or by relying on often unreliable intuition and heuristics.
  • Failure to Use Pretrained Models. It is unclear how to integrate pretrained model components from models with different architectures. Pretrained models are a key advantage of foundation models. However, due to compatibility issues, hybrids are often trained from scratch, leading practitioners to resort to small hybrids in limited settings or incur high costs.
A potential solution to the latter challenge is the use of model merging techniques [42, 44, 41, 15, 8, 16], some of which can operate cross-architecture [1, 12]. Unfortunately, such tools are embryonic–they are expensive and it is unclear how well they work with the diverse types of architectures a user may seek to build a hybrid from.
We propose a framework for automatically designing hybrid architectures that overcomes these obstacles. Our approach is inspired by principles from neural architecture search (NAS), but applies these at the level of LM blocks rather than convolutional cells [23, 21] or operations [36, 35]. The resulting framework is simple and tractable. It sidesteps merging different architectures by using simple linear projectors to translate between the “languages” spoken by various architectures. This enables us to include blocks from many different architectures with little to no changes required. In addition, inspired by the mechanistic architecture design (MAD) framework [34], we show how our framework can learn hybrid architectures via MAD that transfer to new tasks.
Concretely, our proposed system, Manticore:
  1. 1.
    Automatically selects language models, without training several models from scratch,
  2. 2.
    Automatically constructs pretrained hybrids without evaluating the entire search space,
  3. 3.
    Provides a technique for programming hybrids to have certain skills without full training.
Experimentally, our automatically designed hybrids compete with existing hybrids and models on the MAD tasks [34] and Long Range Arena (LRA) [38], we produce pretrained hybrids that can improve downstream fine-tuning performance, and we can program hybrids using the MAD tasks.
2 Related work
Language Model Architectures: Transfomers and Beyond. Transformers are currently the dominant LM architecture. The success of the “vanilla” architecture introduced by Vaswani et. al. [40] has led to many proposed variations. The quadratic complexity of the base self-attention operation has inspired the search for alternative architectures that offer comparable performance with subquadratic complexity. One line of work builds off state-space models, with variations made to enable language modeling [32, 33, 13, 3]. Another line of work involves linear attention by formulating transformers as RNNs and expressing self-attention as a kernel dot-product [17]. Other new approaches increase the expressivity of this formulation with data-dependent gating [43].
Our work does not propose a new architecture. Instead, we focus on the idea that practitioners should be able to take advantage of new architectures in a transparent way.
Neural Architecture Search & Mechanistic Search. Neural architecture search (NAS) techniques are used to automatically search for optimal architectures. These techniques have produced state-of-the-art models in several domains. Much of the challenge in NAS is the complexity of search; in the most standard form, NAS involves a difficult bilevel optimization over a large search space. Much effort has been aimed at reducing costs, often via continuous relaxations of search spaces, with techniques like DARTS [23] and DASH [36].
Using NAS to discover architectures for language modeling—and especially those that may rival Transformers—has thus far been hard. A promising approach is the MAD framework [34] , which uses “mechanistic tasks” (synthetic tasks organized around simple principles) to search for high-quality architectures. While we do not seek to discover new architectures, we are inspired by this approach in our effort to search for hybrid architectures.
Hybrid Architectures. Perhaps unsurprisingly, there is no single dominant architecture among either standards, like Transformers, or emerging subquadratic architectures. While there are some insights that can be converted into heuristics for model selection, generally, to take advantage of new models, practitioners must exhaustively evaluate all of them on each of their tasks. The cost of doing so has inspired the idea of crafting hybrid architectures that mix components from different approaches, with the goal being to obtain best-of-all-worlds behavior.
Unfortunately, the space of hybrid architectures is already large and only grows with each new proposed approach. Manually crafting hybrids is costly; users must either brute-force the enormous search space or alternatively hand-craft a small candidate set of hybrids in the hope that it includes a reasonably performant choice. Our work provides an efficient alternative to this process.
Model Merging. A final prospective approach to using multiple models is merging. Merging pretrained models (of the same architecture) has shown promising results [42, 44, 41, 15, 8, 16], creating powerful large-scale merges such as SOLAR-10.7B [18] and Goliath-120B [^3] https://huggingface.co/alpindale/goliath-120b from two fine-tuned Llama2-70B [39] models. The former two were produced using a trial-and-error-based technique called ‘frankenmerging,’ introduced in MergeKit [12]. Frankenmerging involves stitching together different fine-tuned versions of the same model or, hypothetically, different models. This has inspired efforts to merge models of different architectures using large-scale evolutionary search [1]. However, such efforts are still embryonic, with substantial computational drawbacks, requiring many training runs. Manticore, on the other hand, does not require training a large number of models.
3 Methods
We now describe Manticore, our framework for automatically designing hybrid architectures by mixing components of pretrained models. Manticore relies on projectors to align features across architectures, then applies a convex combination to the aligned features, as summarized in Figure 1. In Section 3.1, we discuss and formally define the structure of Manticore hybrids, including the projectors and convex combination mixture weights, as well as how both of these components are used within Manticore. In Section 3.2, we detail the search procedures (inspired by NAS) and training routines involved in pretraining, fine-tuning, and programming hybrids.
Figure 1: Our proposed Manticore framework, which enables: (1) cross-architecture LM selection, (2) the construction of pretrained hybrids, and (3) the ability to program hybrids to have certain skills.
3.1 The Structure of Manticore Hybrids
Our framework comprises three main parts: the individual LMs that we combine to produce our overall hybrid, projectors that translate feature representations between LMs of different architectures, and convex combination mixture weights that specify how much the hybrid will use the features of each component architecture. We detail each of these in the following.
3.1.1 Component Models
We refer to a model that is used in Manticore as a component model. Any modern LM can be used as a component model in our framework. In this section, we will formally define the general high-level structure of the component models that we support. For an LM $M$ with model embedding dimension $d_{M}$ on a sequence of $t$ tokens from a set $\mathcal{V}$, denoted $x=(x_{1},...,x_{t})\in\mathcal{V}^{t}$, a forward pass $M(x)$ is typically computed using the following recipe:
  1. 1.
    Apply an embedding function, $M_{\text{embed}}:\mathcal{V}^{t}\to\mathbb{R}^{t\times d_{M}}$ to the tokens, resulting in a sequence of embeddings denoted $x_{\text{embed}}=M_{\text{embed}}(x)$.
  2. 2.
    Take forward passes through $L_{M}$ ‘blocks’–we denote the $\ell^{\text{th}}$ block as $M_{\text{Block}}^{(\ell)}:\mathbb{R}^{t\times d_{M}}\to\mathbb{R}^{t\times d_{M}}$. Specifically, for all $\ell\in[L_{M}]$, we obtain $x_{\ell+1}=M_{\text{Block}}^{(\ell)}(x_{\ell})$, where $x_{1}:=x_{\text{embed}}$.
  3. 3.
    Finally, we pass $x_{L_{M}+1}$ into a language modeling head, $M_{\text{head}}:\mathbb{R}^{t\times d_{M}}\to(\Delta^{|\mathcal{V}|-1})^{t}$, where $\Delta^{|\mathcal{V}|-1}$ is the probability simplex of dimension $|\mathcal{V}|$.
This recipe applies to virtually all modern transformer-based LMs, recurrent models, and state-space models. Our framework supports all of these, and any other architecture that follows this recipe.
3.1.2 Projectors
Suppose we have two pretrained component models, $M$ and $M^{\prime}$. In general, even if the model dimensions are the same for both models ($d_{M}=d_{M^{\prime}}$), blocks from $M$ and $M^{\prime}$ may not be directly compatible, as their input and output features are likely to be very different. It is also possible that $d_{M}\neq d_{M^{\prime}}$, in which case composing blocks from $M$ and $M^{\prime}$ is not even well-defined.
Figure 2: Examples of component models used in Manticore. Transformer and Mamba component models are shown.
Figure 3: The projectors with residual connections used in Manticore, used for translating features between pretrained blocks of different component models.
To overcome this issue, we apply projectors to both the inputs and the outputs of a block (or a sequence of blocks, discussed in Section 3.1.4) that we wish to combine in Manticore hybrids. Overall, our goal in designing projectors is to enable the blocks of $M$ and $M^{\prime}$ to speak a common language, such that their features are compatible and can be reused in the resulting hybrid model. This is conceivably challenging—the mapping between feature spaces could be highly nonlinear and might require a lot of task-specific data to adequately learn the mapping. So do projectors need to be heavyweight, data-hungry, highly nonlinear objects? Fortunately, the answer is no—we find that a simple linear transformation with a gated residual, pretrained on general language data, is sufficient.
Suppose that we want to create a Manticore hybrid from $K$ different pretrained component models, denoted $M_{(1)},...,M_{(K)}$ with model dimensions $d_{M_{(1)}},...,d_{M_{(K)}}$. We define $d_{\max}:=\max_{k\in[K]}d_{M_{(k)}}$, then want input and output projectors for the blocks of each model that convert their features to a common feature space of dimension $d_{\max}$. For any sequence of blocks of length $(n+1)<L_{d_{M_{(k)}}}$ from model $M_{(k)}$ and length-$t$ input,
$\left(M_{(k)\text{Block}}^{(\ell+n)}\circ...\circ M_{(k)\text{Block}}^{(\ell)}\right):\mathbb{R}^{t\times d_{M_{(k)}}}\to\mathbb{R}^{t\times d_{M_{(k)}}},$
[…]
$\circ M_{(k)\text{Block}}^{(\ell)}\circ\text{Proj-in}^{(\ell)}_{(k)}\bigg{)}:\mathbb{R}^{t\times d_{\text{max}}}\to\mathbb{R}^{t\times d_{\text{max}}}.$
For input $x\in\mathbb{R}^{t\times d_{M_{(k)}}}$ we parameterize each projector as a linear transformation with gated residual:
[…]
Respectively, $\text{Trunc}(\cdot;d)$ and $\text{Pad}(\cdot;d)$ truncate and zero-pad input to dimension $d$, and $\text{Linear}_{d\to d^{\prime}}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ is a learnable linear transformation. Gating weights are parameterized as $\alpha\in[0,1]$.
In total, where $\alpha\in\Delta^{K-1}$ and $I_{k}$ is a length-$n_{k}$ vector of block indices from component model $k$, we define the output of the block sequence defined by $I_{k}$ as
$h_{k}(x;\alpha_{k},I_{k})=\left(\text{Proj-out}^{(I_{k,n_{k}})}_{(k)}\circ M_{(k)\text{Block}}^{(I_{k,n_{k}})}\circ...\circ M_{(k)\text{Block}}^{(I_{k,1})}\circ\text{Proj-in}^{(I_{k,1})}_{(k)}\right)(x;\alpha_{k}).$
3.1.3 Mixture Weights
Next, we would like to mix the activations of different component models’ block sequences, in a way that allows us to learn how much influence the blocks from each component model will have on the overall hybrid model. Learning the amount of influence that each block sequence should have on the overall hybrid is critical—if certain blocks produce less helpful features, we need a way to down-weight them. Conversely, we want to use the best blocks in our hybrid as much as possible—we want to up-weight these helpful blocks. Overall, a parameterization that allows us to learn these weights should lead to better hybrids.
We do this by taking a convex combination of the projectors’ outputs: given the projected features $h_{k}(x;\alpha_{k},I_{k})$ for each component model $k\in[K]$, we output a convex combination of projected features
$\text{Mix}_{\alpha}(x;I_{1},...,I_{K})=\sum_{k\in[K]}\alpha_{k}h_{k}(x;\alpha_{k},I_{k}).$
Figure 4: Mixture weights used in Manticore, which learn how much influence component model blocks should have.
We reuse the convex combination weights as the gating weights in the projectors. This choice yields the convenient property that when the mixture weights $\alpha$ are set to one in index $k$ and zero everywhere else, the Mix function exactly computes a sequence of blocks from component model $k$ while ignoring the projectors and the blocks from other component models. We adopt a popular parameterization for mixture weights from the NAS literature [23]: we parameterize $\alpha$ using a softmax of scalars. That is, we define $\alpha_{k}:=\frac{\exp(a_{k})}{\sum_{j\in[K]}\exp(a_{j})}$ for all $k\in[K]$.
3.1.4 Manticore
We are now ready to define our overall hybrid architecture. We seek to create a hybrid from $K$ component models, $M_{(1)},...,M_{(K)}$, each with a potentially different number of blocks, denoted $L_{M_{(k)}}$ for component model $k$. We fix $L$ to be the number of Manticore blocks, where $L$ is a common factor of each of the depths $L_{M_{(k)}}$, for all $k\in[K]$—we treat this choice of factor as a hyperparameter. For each of the $L$ Manticore blocks, we want to mix a sequence of blocks from each of the $K$ component models. We also want the number of blocks from each model $k\in[K]$ that are allocated to a single Manticore block to be evenly spread out throughout the $L$ Manticore blocks—this is why we require $L$ to be a factor of $L_{M_{(k)}}$.
For each component model $k\in[K]$, divide the indices of the blocks $[L_{M_{(k)}}]$ evenly into $L$ contiguous parts, denoted as $[L_{M_{(k)}}]=(I_{k,1},...,I_{k,L})$. Then, adopting the notation from our component models, a Manticore block is defined as
$\text{Manticore}_{\text{Block}}^{(\ell)}(\cdot):=\text{Mix}_{\alpha^{(\ell)}}(\cdot;I_{1,\ell},...,I_{K,\ell})$
with $\text{Manticore}_{\text{Block}}^{(\ell)}:\mathbb{R}^{t\times d_{\text{max}}}\to\mathbb{R}^{t\times d_{\text{max}}}$, for each $\ell\in[L]$, and $\alpha^{(\ell)}$ being the mixture weights at $\ell$. Next, we initialize a new set of embedding weights and a new task specific (or language modeling) head, and we can finally illustrate a forward pass with a Manticore hybrid model, denoted using the shorthand notation $\text{Manticore}(\cdot):=\text{Manticore}[M_{(1)},...,M_{(K)}](\cdot)$. Let $x=(x_{1},...,x_{t})\in\mathcal{V}^{t}$ be a sequence of $t$ tokens from a set $\mathcal{V}$. The forward pass is computed as follows:
  1. 1.
    Apply the new embedding function $\text{Manticore}_{\text{embed}}:\mathcal{V}^{t}\to\mathbb{R}^{t\times d_{\text{max}}}$ to the tokens, resulting in a sequence of embeddings denoted $x_{\text{embed}}=\text{Manticore}_{\text{embed}}(x)$.
  2. 2.
    Take forward passes through $L$ Manticore blocks, each with dimension $d_{\text{max}}$, concretely, we compute $x_{\ell+1}:=\text{Manticore}_{\text{Block}}^{(\ell)}(x_{\ell})$, where $x_{1}:=x_{\text{embed}}$.
  3. 3.
    Pass $x_{L_{M}+1}$ into a new task-specific or language modeling head, $\text{Manticore}_{\text{head}}:\mathbb{R}^{t\times d_{M}}\to\mathbb{T}$, where $\mathbb{T}$ is the appropriate output space for the learning task.
In NAS terms, our search space is over the set of $L\ni\ell$ mixture weights $\alpha^{(\ell)}\in\Delta^{K-1}$. However, our search space differs from typical gradient-based NAS techniques in the sense that we do not require discretization to derive a final architecture after we obtain our mixture weights. Typically, NAS would involve selecting a single sequence of component architecture blocks at each of the Manticore blocks, usually by taking the $\operatorname*{arg\,max}$ of the mixture weights. Instead, the mixtures themselves are what characterize Manticore hybrids. Nonetheless, if we were to replace the mixture weights $\alpha^{(\ell)}$ with discrete one-hot vectors, we could derive any of the following: the component model architectures themselves, existing hybrid architectures, and ‘frankenmerged’ models [12].
3.2 How To Use Manticore
With Manticore, we can automatically select language models without training every model in the search space, automatically construct pretrained hybrid architectures without significant trial-and-error, and also program pretrained hybrids without full training. In this section, we will discuss the details of how Manticore can be used in each of these three usage scenarios.
Training hybrids from scratch. Manticore can be used to automatically select LMs without training all of the LMs in the search space. Our selection technique is simple: inspired by gradient-based NAS techniques [23] and treating the mixture weights as our ‘architecture parameters,’ we proceed in two steps: 1. train mixture weights along with all other parameters, and 2. freeze the mixture weights and retrain the rest of the parameters from scratch. Unlike NAS, we found that in many pretraining settings, it was sufficient to stop at 1. and forgo retraining. In our pretraining experiments, we primarily use randomly-initialized GPT-Neo [5] and Mamba [13] as component models without projectors, and separately experiment with a subset of the blocks from MAD [34].
Fine-tuning pretrained hybrids. Manticore can be used to create and fine-tune pretrained hybrids. We create pretrained hybrids as follows: begin with a set of pretrained models, replace their LM heads and embeddings with a single randomly initialized LM head and embedding layer, and pretrain the projectors on a small amount of general language data such as FineWeb [30] while keeping the original component model weights frozen.444 We found one billion tokens to be sufficient for projector pretraining. To fine-tune the pretrained hybrids on downstream task data, we first search for mixture weights by training all of the parameters simultaneously, we freeze the mixture weights, rewind the component models and projectors to their pretrained state, and fine-tune. This procedure completely sidesteps large-scale pretraining of new hybrids. In our synthetic experiments, we create pretrained Manticore hybrids from pretrained GPT-Neo-125M [5] and Mamba-130M [13] models, while for our experiments on real natural language data, we opt for pretrained Pythia-410M [4] and Mamba-370M [13] as component models.
Programming hybrids. Excitingly, we can program Manticore mixture weights by using external information to predict them. We consider two scenarios. If we know that a component model has blocks that are somehow incompatible with the target task—e.g. resulting from sequence length constraints—we can omit these blocks by setting their mixture weights to 0. Otherwise, we can predict good mixture weights by searching on a fixed set of proxy tasks—for this, we use the MAD tasks [34]. The MAD tasks are synthetic unit tests that are predictive of hybrid LM scaling laws, but within our framework, we find that they are also useful for finding general-purpose pretrained hybrids. We use the following procedure for programming mixture weights using the MAD tasks. First, run search on the MAD tasks using a smaller, randomly initialized version of our pretrained hybrid. For each MAD task, our search procedure returns a set of mixture weights—we simply average the resulting mixture weights, freeze them, and fine-tune on the downstream task data.
4 Experimental Results
We provide experimental evidence that validates the following claims about Manticore:
  • C1. Pretrained hybrids can outperform their component models on fine-tuning tasks,
  • C2. Trained from scratch, Manticore hybrids are competitive with existing hybrids and LMs, and
  • C3. We can program mixture weights using external sources without search on the task data.
4.1 Fine-Tuning Pretrained Hybrids
We evaluate C1, first on a synthetic fine-tuning task, and then on natural language fine-tuning tasks.
Setup. We consider a synthetic LM dataset comprising GPT-Neo and Mamba generated completions of text from Penn Treebank [27]. Naturally, we also use pretrained GPT-Neo-125M and Mamba-130M models as component models, creating a single Manticore block with projectors that were pretrained on one billion tokens from FineWeb [30]. We perform search using DARTS and perform post-search retraining with the model weights and projectors rewound to their pretrained state.
Results. Our results are shown in Figure 5 (left). We compare our search results to a sweep over a range of possible mixture weights, and find that our search procedure returns the optimal mixture weights, outperforming both Mamba and GPT-Neo. This confirms our claim that Manticore hybrids can outperform their component models on synthetic fine-tuning tasks. Given that this task comprises two slices that each of our component models should be good at—GPT-Neo should be good at predicting GPT-Neo outputs, and vice versa—we hypothesize that Manticore hybrids are well suited to situations in which the component models have complementary ‘skills’ [7].
Figure 5: Mixture weight sweeps on Penn Treebank completions using pretrained GPT-Neo-125M and Mamba-130M as our component models. (Left) When we create one Manticore block, there is a region of the search space where we improve over Mamba. (Right) The same holds for two Manticore blocks, and our technique for hybrid programming using MAD discovers this region.
Setup. We evaluate on three natural language fine-tuning datasets: Penn Treebank [27], the Alpaca instructions dataset [37], and ELI5 [10]. We use Pythia-410M and Mamba-370M as component models, and create a single Manticore block from their blocks with projectors that were pretrained on one billion tokens from FineWeb [30]. As before, we first search for mixture weights, and then we retrain with the fixed mixture weights found by search.
Results. Our results are shown in Table 1. Manticore outperforms its component models on Alpaca and ELI5, while it achieves performance between its two component models on Penn Treebank. This confirms our claim that Manticore can outperform component models on real natural language tasks. The fact that Mamba-370M outperforms Manticore in this setting is not a failure of our framework, as Mamba-370M is included as part of our search space. We speculate that the use of more powerful search procedures from the NAS literature, such as GAEA [21], could improve our search performance and help to recover or outperform Mamba-370M.
Table 1: Manticore on natural language tasks using Pythia-410m and Mamba-370m as component models. The best test losses are bolded and the second-best are underlined.
Task Pythia-410M (A) Mamba-370M (B) Manticore [A, B]
Penn Treebank 0.9099 0.8397 0.8600
Alpaca 2.5011 2.2999 2.1779
ELI5 4.1260 3.9414 3.9331
4.2 Training Hybrids from Scratch
For C2, we compare to prior hybrids on MAD and non-hybrid component models on LRA and MAD.
Setup. We compare training Manticore from scratch to training existing hybrid architectures on the MAD tasks. We begin with two hybrid architectures from the literature: Mambaformer [29], which combines Mamba and attention blocks, and the striped multi-head Hyena + Mixture-of-Experts (MoE) MLP architecture that was shown to perform well on the MAD tasks [34]. We compare these two baselines to a Manticore hybrid combining three component models: striped multi-head Hyena + MoE-MLP, a transformer, and Mamba. We use two blocks for each of these architectures, creating two Manticore blocks. Again, we search for mixture weights and then retrain.
Results. The results of this experiment are shown in Table 2. We outperform the striped multi-head Hyena + MoE model from the MAD paper, and we approach the performance of Mambaformer on all but one task. This validates the claim that Manticore hybrids, trained from scratch, compete with existing hybrids. Despite Mambaformer not being a component model, it is in our search space, and we again speculate that improvements in search would lead to its recovery.
Table 2: Trained from scratch on MAD tasks, Manticore beats or matches the performance of existing hybrids on all but one task. The best test losses are bolded and the second best are underlined.
Task Striped MH Hyena Mambaformer Manticore
+ MoE-MLP
In-context Recall 3.7153 0.0020 0.0048
Fuzzy In-context Recall 4.1714 4.1712 4.1750
Noisy In-context Recall 4.1643 4.1646 4.1607
Selective Copying 1.8021 0.0005 0.0171
Memorization 8.8353 5.2179 8.9254
Setup. We compare Manticore hybrids to their component models on LRA, when trained from scratch. We create GPT-Neo and Mamba component models of similar sizes to those in Tay et al. [38] and create a Manticore hybrid. As a simplified pipeline, we do not retrain model weights after search.
Results. Our results are shown in Table 3. We outperform the component models on all tasks except for IMDb, in which case Manticore was between GPT-Neo and Mamba. This validates the claim that Manticore hybrids, trained from scratch, compete with existing LMs.
Table 3: Manticore hybrids trained from scratch on LRA using GPT-Neo and Mamba component models. The best test accuracies are bolded. GPT-Neo does not support the Pathfinder-X sequence length requirement, so we set its mixture weight to 0 and Manticore reduces to Mamba.
Task GPT-Neo (A) Mamba (B) Manticore [A, B]
ListOps 37.90 20.65 38.70
IMDb 59.62 87.74 72.44
CIFAR10 39.37 20.81 43.15
Pathfinder32 89.41 85.76 91.45
Pathfinder-X N/A $\textbf{75.50}^{*}$ $\textbf{75.50}^{*}$
Setup. Next, we compare Manticore to non-hybrid architectures trained from scratch on the MAD tasks. We compare two-block GPT-Neo and Mamba models to a Manticore hybrid using a single Manticore block. Again, we report the performance of the search procedure itself without retraining.
Results. Our results are shown in Table 4. Manticore outperforms GPT-Neo and Mamba on all of the MAD tasks in this setting. This provides further evidence for our claim that Manticore hybrids compete with existing LMs when trained from scratch. It is conceivable that our larger Manticore hybrids simply perform better than component models due to their size—however, we find that post-search discretization and retraining tends to result in similar performance, but reduces the model size by roughly half. We include an ablation of post-search discretization in the Appendix.
Table 4: Trained from scratch on the MAD tasks, Manticore improves over small two-block component models combined into a single Manticore block. The best test losses are shown in bold.
Task GPT-Neo (A) Mamba (B) Manticore [A, B]
In-context Recall 4.0771 4.1858 4.0768
Fuzzy In-context Recall 4.4384 4.8097 4.2797
Noisy In-context Recall 4.1843 4.2605 4.1823
Selective Copying 1.0470 3.7765 0.9478
Memorization 4.6110 5.2281 4.1367
4.3 Programming Hybrids
We evaluate C3 with two types of external data: access to task metadata such as sequence length requirements, and the use of the MAD tasks as a proxy for search on downstream task data.
Setup. As in many of our previous experiments, we used the GPT-Neo and Mamba architectures as component models to our Manticore hybrid. However, this time, we set out to train from scratch on the extremely long-range Pathfinder-X task from LRA, which requires sequence length support greater than that of GPT-Neo. Using this external information about the task, we set the mixture weights for GPT-Neo to 0, which in this case, means that Manticore reduces to Mamba. [^5] As of writing, Mamba on LRA is open: https://github.com/state-spaces/mamba/issues/282.
Results. The results of this experiment are shown in the last row of Table 3. In the simple case of having access to task metadata, this validates the claim that we can program mixture weights to exclude incompatible blocks. At the time of writing, we are not aware of prior published Mamba results on LRA despite community interest, which would make our evaluation in Table 3 the first such result. Note that we did not thoroughly tune hyperparameters, so we view this result as a starting point for the community, rather than a final answer.
Setup. Finally, in the case in which we can run all of our component models on our learning task, we program the mixture weights using the MAD tasks as a search proxy. We set out to fine-tune a pretrained hybrid comprising GPT-Neo-125M and Mamba-130M with two Manticore blocks on our Penn Treebank completions synthetic. We train a scaled-down version of this Manticore hybrid with randomly initialized weights and two blocks per component model on the MAD tasks. This yields mixture weights for each of the MAD tasks—we average them and then fine-tune our pretrained hybrid on Penn Treebank completions using the predicted mixture weights.
Results. Our results are shown in Figure 5 (right). We superimpose the predicted mixture weights and mean search trajectory from MAD onto the architecture loss landscape computed on Penn Treebank completions. We find that this procedure recovers a hybrid that outperforms the component models (Mamba, lower right; GPT-Neo, upper left) and substantially outperforms the naive frankenmerges in our search space (upper right and lower left) [12]. This validates the claim that we can program mixture weights using external sources without performing search on the task data. Intriguingly, search on the MAD tasks appears to follow the architecture gradient on the different downstream fine-tuning task, even though the architecture is scaled-down and trained from scratch on MAD. We suspect that the mixture weights and architecture loss landscapes for pretrained hybrids are fairly universal across fine-tuning tasks, and that the same procedure is likely to work on a wide range of scenarios. Furthermore, we hypothesize that this technique could outperform other gradient-based NAS methods directly applied to the downstream task.
Appendix
Appendix A Ablations
Choice of search algorithm. By default, we use a form of the single-level DARTS [23] search algorithm in all of our experiments requiring search. We optionally evaluate whether or not to take alternating update, that is, we alternately take gradient steps in the architecture and model parameters—we treat this choice as a task-dependent hyperparameter. However, there are many alternative NAS algorithms that we could have used for search. In our ablation of the choice of search algorithm, we also evaluate DASH [36] on our Penn Treebank completions synthetic—the results of which are shown in Table 5. In general, we found that using DASH was unable to recover strong architectures in our search space. We postulate that this is because DASH simply aims to solve a different problem, and is not suited to our search space: namely, DASH is used to search for lower-level operations, rather than LM blocks. We also found that alternating DARTS updates was somewhat helpful, compared to simultaneously updating all of the parameters at once—for our experiments, we treated this choice as a hyperparameter.
Table 5: Comparison of NAS search methods on our Penn Treebank completions synthetic.
Alternating? DARTS DASH
Yes 1.2854 2.5899
No 1.3635 2.5968
Whether or not to discretize after search. We perform an ablation of whether or not to perform discretization on our MAD task experiments in which we compare to existing hybrids. We find that while discretization can sometimes improve performance, the performance differences are often marginal. If final parameter count is a concern, then discretization is beneficial.
Table 6: A comparison of non-discretized vs. discretized Manticore.
Task Manticore Manticore
(non-discretized) (discretized)
In-context Recall 0.0068 0.0081
Fuzzy In-context Recall 4.1764 4.1729
Noisy In-context Recall 4.1628 4.1614
Selective Copying 0.0849 0.0006
Memorization 8.9416 8.9402
Appendix B Additional MAD results
In the main text of the paper, we presented results comparing Manticore hybrids trained from scratch to existing hybrids from the literature—namely Mambaformer and the Striped MH Hyena + MOE architecture from MAD. Notably, the Striped MH Hyena + MOE architecture was only the second best architecture presented in the MAD paper. We found that their best architecture, the Striped Hyena Experts + MOE model, performed slightly worse on the harder versions of the MAD tasks that we evaluated. We present these results in Table 7.
Table 7: Trained from scratch on MAD tasks, Manticore beats or matches the performance of existing hybrids on all but one task. The best test losses are bolded and the second best are underlined.
Task Striped Hyena Experts Striped MH Hyena Mambaformer Manticore
+ MoE-MLP + MoE-MLP
In-context Recall 4.0315 3.7153 0.0020 0.0048
Fuzzy In-context Recall 4.1749 4.1714 4.1712 4.1750
Noisy In-context Recall 4.1640 4.1643 4.1646 4.1607
Selective Copying 2.1731 1.8021 0.0005 0.0171
Memorization 8.8537 8.8353 5.2179 8.9254
Appendix C Additional Pathfinder Results
We ran several additional variants of the pathfinder task for which the required sequence length exceeded the maximum supported sequence length of GPT-Neo. We report these results in Table 8.
Table 8: Additional Pathfinder results. Note that since these variants of Pathfinder exceed the maximum sequence length of GPT-Neo, we set its mixture weight to be 0 and evaluate using Mamba.
Pathfinder task GPT-Neo Mamba Manticore
(A) (B) [A, B]
$64\times 64$, 6 paddles N/A 80.40 80.40
$64\times 64$, 9 paddles N/A 90.01 90.01
$64\times 64$, 14 paddles N/A 86.87 86.87
$128\times 128$, 6 paddles N/A 75.50 75.50
Appendix D Hyperparameters
D.1 Fine-Tuning Pretrained Hybrids
Penn Treebank completions synthetic. For model weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5. For mixture weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $0.005$ and use alternating updates.
Fine-tuning on language tasks. For model weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5. For mixture weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $0.005$ and use simultaneous updates.
D.2 Training Hybrids from Scratch
Comparison to existing hybrids on MAD.
We provide the hyperparameters and training details for our MAD evaluations from Section 4.2
Existing hybrids were trained with a hyperparameter grid search over the space $[1e-4,5e-4,1e-3]$ for learning rate and $[0.0,0.1]$ for weight decay, similar to the procedure in MAD [34].
Manticore is trained in two stages. In the first stage, we train the model and architecture weights in the alternating schedule utilized in DARTS [23]. In this stage, we perform a hyperparameter grid search of the space $[1e-4,5e-4,1e-3]$ for model weight learning rate, $[1e-4,1e-4]$ for architecture weight learning rate, and $[0.1]$ for weight decay. In the second stage, the architecture weights are frozen and we train only the model weights using the best learning rate found in the first stage.
Evaluation on LRA. We provide the hyperparameters and training details for our LRA evaluations.
  • ListOps. We trained all models with 5000 steps. The hyperparameter for GPT-Neo is 8 heads, 6 layers, 512 as the embedding dimension, and 2048 as FFN dimension. The hyperparameter for Mamba is 12 layers, with 512 as the model dimension. The vocab size is 18.
  • IMDb. We trained all models with 25 epochs and batch size 32.The hyperparameter for GPT-Neo is 8 heads, 6 layers, 512 as the embedding dimension, and 2048 as FFN dimension. The hyperparameter for Mamba is 12 layers, with 512 as the model dimension. The vocab size is 129.
  • CIFAR10. We trained all models with 10 epochs. The hyperparameter for GPT-Neo is 4 heads, 3 layers, 64 as the embedding dimension, and 128 as FFN dimension. The hyperparameter for Mamba is 6 layers, with 64 as the model dimension. The vocab size is 256, which is the pixel value range of the grayscale image.
  • Pathfinder32. We trained all models with 10 epochs. The hyperparameter for GPT-Neo is 8 heads, 4 layers, 128 as the embedding dimension, and 128 as FFN dimension. The hyperparameter for Mamba is 8 layers, with 128 as the model dimension. The vocab size is 256, which is the pixel value range of the grayscale image.
Comparison to non-hybrids on MAD.
We use two blocks each from GPT-Neo and Mamba, each with a model dimension of 128. We train for 200 epochs and select the best performance during training, as all of the models overfit across the board. We use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5.
D.3 Programming Hybrids
Mamba evaluation on long Pathfinder tasks. Due to our limited computation resources, we did not conduct a hyperparameter sweep for the result we presented. We used Mamba with models of a similar size as Pathfinder32, which has 8 layers, 128 as the hidden dimension size, and 256 as the vocab size. The $64\times 64$, 6 paddles version is trained by 10 Epoch with default HP. The result for other versions is trained with 200 epochs with default HP in Huggingface trainer.
MAD tasks as a search proxy. For model weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5. For mixture weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $0.01$ and use simultaneous updates. For search on the MAD tasks, we train scaled-down versions of GPT-Neo and Mamba each with four blocks, model dimensions of 128, and no projectors.
D.4 Pretraining Projectors
For all non-frozen weights (i.e., projectors, mixture weights, embeddings, and the LM head), we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5.
Appendix E Data and MAD Task Parameters
We provide a more detailed description of the datasets that we use in our experiments. We perform our experiments on a range of synthetic and real tasks that measure various aspects of modern LM capabilities. We discuss the specific datasets that we use in our experiments below. MAD synthetics. The MAD synthetic datasets are a set of tasks introduced by Poli et al. [34] to systematically evaluate the design space of LMs. These tasks are designed to serve as proxy unit tests for rapidly prototyping of new hybrid LM architectures. In our experiments, we use harder variants of the MAD tasks, in which we use a larger vocabulary size of 128 instead of the default 16 for most of the tasks, along with fewer training examples. For simplicity, we omit the compression task as it requires the use of encoder-decoder architectures.
  • In-context recall. MAD utilizes a multi-query associative recall task, challenging models to retrieve values linked to keys within input sequences, testing their in-context learning ability across randomly shuffled mappings. We use a vocab size of 128 and 800 training examples.
  • Fuzzy in-context recall. This is a variant of in-context recall to assess a model’s ability to semantically group adjacent tokens. Variable-length keys and values are randomly paired, testing the model’s capacity for fuzzy recall. We use a vocab size of 128 and 800 training examples.
  • Noisy in-context recall. This is an adaptation of in-context recall to evaluate a model’s capacity to disregard irrelevant information. This involves inserting tokens from a separate vocabulary randomly among key-value pairs, enhancing the memorization challenge. We use a vocab size of 128, a noise vocab size of 16 with 80% noise, and 800 training examples.
  • Selective Copying. MAD employs a selective copying task to evaluate a model’s ability to remember and replicate specific tokens from an input sequence while disregarding randomly inserted noise tokens, emphasizing the preservation of token order. We use a vocab size of 128 with 96 tokens to copy, and 800 training examples.
  • Memorization. MAD assesses language models’ factual knowledge retention through a memorization task, where models learn fixed key-value mappings without in-context computation, testing pure memorization ability. For this task, we use a vocab size of 8192.
Long Range Arena. Long Range Arena (LRA) [38] is a benchmark consisting of various tasks of different modalities that evaluate how well models can learn long-context data. For simplicity, we omit byte-level document retrieval as it requires two forward passes per example.
  • Long ListOps. This task is designed to understand whether the architecture is able to model hierarchically structured data in a long-context [28].
  • Byte-level text classification. This task attempts to test the model’s ability to deal with compositionality as in the real world, the model needs to compose characters into words and words into higher-phrases in not so well defined boundaries making it a challenging task, we use IMDB dataset[25] in the LRA paper [38].
  • Image classification on a sequence of pixels. This task aims to understand whether a model is able to capture the 2D spatial structure when presented with a flattened 1D version of an image to classify, we use pixel information from CIFAR10[20] dataset.
  • Pathfinder. This task helps to understand whether a model can reason about whether the given 2 dots in an image are connected by a path having dashes or not. The sequence length is 1024 i.e a 32x32 image is flattened and provided as input to the model [22, 19].
  • Pathfinder-X. An extreme version of Pathfinder with a higher resolution, such as 64x64 and 128*128, which results in a sequence length of up to 16K
Penn Treebank completions. We generate a synthetic dataset of generated text from pretrained GPT-Neo-125M [5] and pretrained Mamba-130M models [13]. We prompt both models using the first four words of every example in the Penn Treebank [27] validation set, which yields two natural slices of our dataset: sentence completions generated by GPT-Neo and those generated by Mamba.
Natural language tasks. We evaluate the ability to fine-tune Manticore on natural language datasets. Specifically, we evaluate on Penn Treebank [26], the Alpaca instruction tuning dataset [37], and an i.i.d. split of the ELI5 training set [10]. Additionally, we use one billion tokens sampled from the FineWeb dataset [30] to pretrain our projector weights.
Appendix F A Call for Action & Community Recommendations
Throughout our research process, we noted a handful of opportunities that help to democratize LM research. Should these opportunities be taken up by the research community, we believe they could help to democratize and help to decentralize community-driven LM research, all which enabling further research on pretrained hybrids.
A search engine for pretrained models.
Surprisingly, we were unable to easily search for pretrained LMs of certain sizes or with certain properties (using Huggingface or otherwise). Tools like this should exist: this would not only significantly democratize LMs, but it would help to reduce monopolies on LM releases and usage, and thereby decentralize LM research.
Standardized, block-structured LM implementations.
We found that standard tools such as Huggingface and PyTorch were insufficient to cleanly access intermediate activations across several model implementations. This could be resolved by adopting standard implementations or structures for LMs that share the common block structure that we describe in Section 3.1.1. Instead, our solution was to fork implementations of several Huggingface models, which is time-consuming, error-prone, and non-scalable. A solution to this problem would enable and encourage further research on pretrained hybrid models, which in turn helps to democratize LM research.
Removing tokenizers from LM pipelines.
We believe that there are too many possible tokenizers, and that tokenizers have a significant potential to introduce merge conflicts in model merging/pretrained hybrid pipelines. In response to this challenge, in our work, we simply chose an arbitrary tokenizer and relearned our embeddings and LM head from scratch in all of our experiments. Possible solutions to this problem would be: as a community, we agree on a standard (small) set of tokenizers, or we eliminate tokenizers altogether by learning character or byte-level LMs.
Appendix G Limitations
At various points in Section 4, we described limitations with using DARTS (the off the shelf NAS search algorithm that we used) for search, in that it was not always able to recover the best architecture in the search space. A potential limitation of Manticore is that it relies on the existence of good gradient-based NAS search algorithms, potentially tailored to our search space. However, we postulate that this is possible, and we leave the task of developing new search techniques to future work.
Appendix H Compute Resources
We ran our experiments on the following GPU hardware:
  • 2x Nvidia RTX A6000 GPUs with 48GB GPU memory hosted locally in a nook in the lead author’s house and in a friend’s basement.
  • 2x Nvidia RTX 4090 GPUs with 24GB GPU memory each hosted locally in other friends’ basements.
  • 2x Nvidia Tesla V100 GPUs with 16GB GPU memory each hosted on AWS (p3.2xlarge instances).
In total, we estimate that our total number of GPU hours across all experiments (those which failed as well as those included in the paper) amounted to roughly 750 GPU hours. We estimate that less than half of these hours accounted for experiments that were not ultimately included in the paper.
Appendix I Broader Impacts and Safeguards
We acknowledge the possibility of misuse with respect to any form of LM research. In our work, among other things, we enable the creation of pretrained hybrid models from existing pretrained models. This has potentially positive and negative social impacts for the community. As a positive potential social impact, we enable the community to much more easily create their own hybrid models of various sizes without large scale pretraining—this has as much potential for positive impact in that these models can be used for good. On the other hand, the ability to create large pretrained hybrids, potentially with custom sets of skills, has the potential to open the door to misuse. To safeguard against such things, we will include appropriate licenses and rules for usage when we ultimately deploy a Python package for the community to more broadly use our framework.
Appendix J Expanded Version of Figure 5 (Right)
To show how the architectures evolve over search on all of the MAD tasks in our mixture weights programming experiment, we provide a more detailed version of Figure 5 (Right) – this is shown in Figure 6. Here, we plot the architecture trajectories throughout training on all of the MAD tasks, and superimpose them onto the architecture-loss landscape of the Penn Treebank completions task. The trajectories roughly follow what appears to be a gradient in the loss landscape, and all of the trajectories are roughly similar. We derive our final ‘programmed’ alphas by taking the average of the final alpha values on each of the MAD tasks, after training.
Figure 6: Mixture weight sweeps on Penn Treebank completions using pretrained GPT-Neo-125M and Mamba-130M as our component models. There is a region of the search space where we improve over Mamba when using two Manticore blocks, and our technique for hybrid programming using MAD discovers this region.